# Evaluating BM3D and NBNet: A Comprehensive Study of Image Denoising Across Multiple Datasets

### v1

@@ Abstract
Index Terms:
Image denoising, non-local filtering, SHAP analysis, image quality assessmentI Introduction
In the field of image denoising, a multitude of techniques have been developed and refined over the years [1, 2]. In this study, we concentrate on two well-established categories of denoising methods: traditional non learning-based techniques, and more recent learning-based approaches. Traditional methods, such as Non-Local Means (NLM) [3] and Block-Matching 3D (BM3D) [4], leverage local similarities within an image and the independent nature of noise. BM3D, one of the methods that we used here, works by searching for similar 2D image blocks within an image and stacking them in 3D arrays. This process takes advantage of the high level of similarity and redundancy in multidimensional data to effectively denoise images. However, this method is very resource consuming[5] and the run-time exponentially grows with the input image size and quality. Another problem with traditional methods such as BM3D is they are not generalizable and for each noisy image or a set of similar noisy images, we need to tune model parameters to achieve the best results. More recent learning-based methods such as deep neural network (DNN)-based denoising strategies, utilize the implicit knowledge of image priors and noise characteristics, acquired through training on paired datasets[6]. While these convolutional neural network (CNN) approaches have seen substantial breakthroughs, they often struggle in complex scenarios, such as in areas with subtle textures or intricate high-frequency details. CNNs typically rely on local filters to differentiate between noise and actual signal. However, with a low signal-to-noise ratio (SNR), these local responses can become unreliable. In our study, we specifically investigate these effects by using the CURE-OR and CURE-TSR datasets which provide multiple levels of the same noise on images. Levels 3 -5 of each challenge in these datasets present a very low SNR. NBNEt[7], the other method that we used in our study, addresses this challenge by incorporating non-local image data through a technique called image projection. The network generates a set of basis vectors from the input image and reconstructs the image within a subspace defined by the span of these vectors. Given that natural images often exist within a low-rank signal subspace, reconstructing the image from this subspace should effectively suppress the noise. NBNet currently holds a state-of-the-art on some image-denoising benchmark. However, the power of both methods remains to be evaluated in this study.
@@ II Datasets
CURE-OR Dataset[8]
@@
CURE-TSR Dataset:[9]
@@
Chest-Xray: [11]
@@ III Methodology
This method operates in two identical steps: initially generating a basic estimate of the noisy image using hard thresholding, followed by the actual denoising using the Wiener filter in the second step. For this purpose, BM3D uses the basic estimate from the first step as a reference or a pilot signal in the Wiener filter [12]. BM3D struggles to generalize across different noise types, and its settings aren’t automatically tailored to the input image’s noise level. To optimize its performance, we used grid-search parameter tuning, where we tested various parameters on a subset of each dataset, challenges, and noise levels and selected the best ones based on the PSNR values of the denoised images. We then applied these optimal parameters to denoise the full set of images.
We also used MATLAB’s parfor [13] function to denoise images in parallel, randomly picking a certain number of images from each set. This was especially important for processing resource-intensive datasets like CURE-OR and CURE-TSR, which have various noise levels, and for large images like those in the SIDD dataset, since BM3D’s computational load depends on the input image size.
@@
In this study, we assessed the performance of the denoising methods across all datasets using seven Image Quality Assessment (IQA) metrics: PSNR, SSIM, CW_SSIM, UNIQUE[14], MS_UNIQUE[15], CSV[16], and SUMMER[17]. Additionally, for datasets with applications in detection or recognition, we incorporated a machine learning-based metric. This comprehensive approach allowed us to evaluate the denoising effectiveness not only through traditional IQA methods but also through a machine-learning perspective on image quality and see how image quality impacts machine learning performance.
@@ IV Discussion
For the CURE-OR dataset, we’ve added an extra evaluation from a machine learning standpoint. We utilized a Faster RCNN architecture, as outlined in Ren et al. (2015)[18], with weights pre-trained on the COCO object dataset [19]. We used the original images as a baseline for reference. Then, we tested the model on images denoised by both methods, measuring performance using an F-1 score derived from precision and recall. This score compares the original and denoised images object detection. Due to the range of noise challenges and the very low SNR in levels 3 to 5 of the CURE-OR dataset, detection performance is notably low—almost impossible in some cases. Training an object detection model from scratch on these noisy images for a sufficient number of epochs could yield much better results by learning the relevant noise distributions. However, given the scope and time constraints of this project, we haven’t been able to train the model on all subsets of this dataset. The methods show varying performances across the different challenges in this dataset which we discussed one by one in the following.
@@
Looking at figure 4 we saw that there is a strong correlation between some of these values thus We employed SHAP analysis [20] utilizing the DeepExplainer from the SHAP library to compute values reflecting each feature’s contribution. Feature importance was gauged through mean absolute SHAP values across the test datasets.
@@ Conclusion
We have followed the instructions for 4 pages + 1 page reference.

### v2

@@ Abstract
I Introduction
In the field of image denoising, a multitude of techniques have been developed and refined over the years [fan2019brief, goyal2020image]. In this study, we concentrate on two well-established categories of denoising methods: traditional non learning-based techniques, and more recent learning-based approaches. Traditional methods, such as Non-Local Means (NLM) [liu2008robust] and Block-Matching 3D (BM3D) [dabov2007image], leverage local similarities within an image and the independent nature of noise. BM3D, one of the methods that we used here, works by searching for similar 2D image blocks within an image and stacking them in 3D arrays. This process takes advantage of the high level of similarity and redundancy in multidimensional data to effectively denoise images. However, this method is very resource consuming[li2017improved] and the run-time exponentially grows with the input image size and quality. Another problem with traditional methods such as BM3D is they are not generalizable and for each noisy image or a set of similar noisy images, we need to tune model parameters to achieve the best results. More recent learning-based methods such as deep neural network (DNN)-based denoising strategies, utilize the implicit knowledge of image priors and noise characteristics, acquired through training on paired datasets[zhang2017beyond]. While these convolutional neural network (CNN) approaches have seen substantial breakthroughs, they often struggle in complex scenarios, such as in areas with subtle textures or intricate high-frequency details. CNNs typically rely on local filters to differentiate between noise and actual signal. However, with a low signal-to-noise ratio (SNR), these local responses can become unreliable. In our study, we specifically investigate these effects by using the CURE-OR and CURE-TSR datasets which provide multiple levels of the same noise on images. Levels 3 -5 of each challenge in these datasets present a very low SNR. NBNEt[cheng2021nbnet], the other method that we used in our study, addresses this challenge by incorporating non-local image data through a technique called image projection. The network generates a set of basis vectors from the input image and reconstructs the image within a subspace defined by the span of these vectors. Given that natural images often exist within a low-rank signal subspace, reconstructing the image from this subspace should effectively suppress the noise. NBNet currently holds a state-of-the-art on some image-denoising benchmark. However, the power of both methods remains to be evaluated in this study.
@@ II Datasets
CURE-OR Dataset[temel2018cure]
@@
CURE-TSR Dataset:[Temel2017_NIPSW]
@@
Chest-Xray: [kermany2018identifying]
@@ III Methodology
This method operates in two identical steps: initially generating a basic estimate of the noisy image using hard thresholding, followed by the actual denoising using the Wiener filter in the second step. For this purpose, BM3D uses the basic estimate from the first step as a reference or a pilot signal in the Wiener filter [hasan2018improved]. BM3D struggles to generalize across different noise types, and its settings aren’t automatically tailored to the input image’s noise level. To optimize its performance, we used grid-search parameter tuning, where we tested various parameters on a subset of each dataset, challenges, and noise levels and selected the best ones based on the PSNR values of the denoised images. We then applied these optimal parameters to denoise the full set of images.
We also used MATLAB’s parfor [MATLAB] function to denoise images in parallel, randomly picking a certain number of images from each set. This was especially important for processing resource-intensive datasets like CURE-OR and CURE-TSR, which have various noise levels, and for large images like those in the SIDD dataset, since BM3D’s computational load depends on the input image size.
@@
In this study, we assessed the performance of the denoising methods across all datasets using seven Image Quality Assessment (IQA) metrics: PSNR, SSIM, CW_SSIM, UNIQUE[temel2016unique], MS_UNIQUE[prabhushankar2018ms], CSV[temel2016csv], and SUMMER[temel2019perceptual]. Additionally, for datasets with applications in detection or recognition, we incorporated a machine learning-based metric. This comprehensive approach allowed us to evaluate the denoising effectiveness not only through traditional IQA methods but also through a machine-learning perspective on image quality and see how image quality impacts machine learning performance.
@@ IV Discussion
For the CURE-OR dataset, we’ve added an extra evaluation from a machine learning standpoint. We utilized a Faster RCNN architecture, as outlined in Ren et al. (2015)[ren2015faster], with weights pre-trained on the COCO object dataset [lin2014microsoft]. We used the original images as a baseline for reference. Then, we tested the model on images denoised by both methods, measuring performance using an F-1 score derived from precision and recall. This score compares the original and denoised images object detection. Due to the range of noise challenges and the very low SNR in levels 3 to 5 of the CURE-OR dataset, detection performance is notably low—almost impossible in some cases. Training an object detection model from scratch on these noisy images for a sufficient number of epochs could yield much better results by learning the relevant noise distributions. However, given the scope and time constraints of this project, we haven’t been able to train the model on all subsets of this dataset. The methods show varying performances across the different challenges in this dataset which we discussed one by one in the following.
@@
Looking at figure 4 we saw that there is a strong correlation between some of these values thus We employed SHAP analysis [cohen2005feature] utilizing the DeepExplainer from the SHAP library to compute values reflecting each feature’s contribution. Feature importance was gauged through mean absolute SHAP values across the test datasets.
@@ Conclusion
Appendix A
Method PSNR SSIM CW_SSIM UNIQUE MS_UNIQUE CSV SUMMER NBNET Blur-1 Blur-2 Blur-3 Blur-4 Blur-5 CodecError-1 CodecError-2 CodecError-3 CodecError-4 CodecError-5 Darkening-1 Darkening-2 Darkening-3 Darkening-4 Darkening-5 Decolorizatio-1 Decolorizatio-2 Decolorizatio-3 Decolorizatio-4 Decolorizatio-5 DirtyLens-1 DirtyLens-2 DirtyLens-3 DirtyLens-4 DirtyLens-5 Exposure-1 Exposure-2 Exposure-3 Exposure-4 Exposure-5 Haze-1 Haze-2 Haze-3 Haze-4 Haze-5 LensBlur-1 LensBlur-2 LensBlur-3 LensBlur-4 LensBlur-5 Noise-1 Noise-2 Noise-3 Noise-4 Noise-5 Rain-1 Rain-2 Rain-3 Rain-4 Rain-5 Shadow-1 Shadow-2 Shadow-3 Shadow-4 Shadow-5 Snow-1 Snow-2 Snow-3 Snow-4
Method PSNR SSIM CW_SSIM UNIQUE MS_UNIQUE CSV SUMMER BM3D Blur-1 Blur-2 Blur-3 Blur-4 Blur-5 CodecError-1 CodecError-2 CodecError-3 CodecError-4 CodecError-5 Darkening-1 Darkening-2 Darkening-3 Darkening-4 Darkening-5 Decolorization-1 Decolorization-2 Decolorization-3 Decolorization-4 Decolorization-5 DirtyLens-1 DirtyLens-2 DirtyLens-3 DirtyLens-4 DirtyLens-5 Exposure-1 Exposure-2 Exposure-3 Exposure-4 Exposure-5 Haze-1 Haze-2 Haze-3 Haze-4 Haze-5 LensBlur-1 LensBlur-2 LensBlur-3 LensBlur-4 LensBlur-5 Noise-1 Noise-2 Noise-3 Noise-4 Noise-5 Rain-1 Rain-2 Rain-3 Rain-4 Rain-5 Shadow-1 Shadow-2 Shadow-3 Shadow-4 Shadow-5 Snow-1 Snow-2 Snow-3 Snow-4